OpenStudy (anonymous):

Please help :) Express the following function, F(x), as a composition of two functions, f and g. f(g(x)). F(x)= (sqrt3x-4) Part 1: Find g(x) by identifying the function you are taking the square root of. Part 2: Find f(x) by identifying what you are doing to the function you found in part 1. What function is f(X)? Part 3: Verify that f(g(x))=f(x) f(g(x))=(sqrt3x-4)

OpenStudy (zzr0ck3r):

ok do you know what composition is?

OpenStudy (anonymous):

yep! but i'm a little confused especially on verification

OpenStudy (zzr0ck3r):

suppose we have the function \(f(x) = \sqrt{x}\) Then \(f(4) = \sqrt{4}=2\) and \(f(25) = \sqrt{25} = 5\) and \(f(3) = \sqrt{3}\) and \(f(2x)=\sqrt{2x}\)

OpenStudy (zzr0ck3r):

what do we need \(g(x) \) to be so that we have \(f(g(x)) = \sqrt{3x-4}\)?

OpenStudy (anonymous):

sqrtx?

OpenStudy (zzr0ck3r):

we are letting \(f(x) = \sqrt{x}\).

OpenStudy (zzr0ck3r):

hint: \(f(\color{blue}{g(x)})= \sqrt{\color{blue}{3x-4}}\)

OpenStudy (zzr0ck3r):

what is \(g(x) = \)?

OpenStudy (anonymous):

oh! √x ?

OpenStudy (zzr0ck3r):

no

OpenStudy (zzr0ck3r):

you are just saying the same thing over and over...

OpenStudy (anonymous):

wait sorry

OpenStudy (anonymous):

i wasn't paying attention

OpenStudy (zzr0ck3r):

right

OpenStudy (anonymous):

3x-4

OpenStudy (zzr0ck3r):

correct

OpenStudy (zzr0ck3r):

ok, so do you see why we choose \(f(x) = \sqrt{x}\)?

OpenStudy (anonymous):

yeah i get that now haha

OpenStudy (zzr0ck3r):

the last part makes no sense. It is not true that \(f(g(x))= f(x)\)

OpenStudy (anonymous):

it then says f(gx)) = (sqrt3x-4) though so is it still totally wrong